# Discord-type quantum correlations in axially symmetric spin-(1/2, $S$) systems

### v1

@@ V.1 Spin Hamiltonian and Gibbs density matrix
The energy levels of the system are obviously $E_{0}$, $E_{4S+1}$ and $2S$ additional pairs
@@
$R_{k}=\sqrt{(h_{k}-h_{2S+k})^{2}+4|g_{k}|^{2}}$ (39)
and $k=1,\ldots,2S$.
The partition function $Z=\sum_{n}\exp(-E_{n}/T)$ of the system is given by
@@
The Gibbs density matrix is equal to
$\rho=\frac{1}{Z}\exp(-{\cal H}/T).$ (42)
Since $\cal H$ has the $\cal A$ form, this functional relationship also leads to the density matrix of the same structure, Eq. (3). If the dimension of the matrix $\cal H$ is finite, $\rho$ can be found manually or obtained on a digital machine by direct symbolic (analytical) calculations. After this, the results can be extended by induction to systems with arbitrary spin $S$. This allows us to get all non-zero matrix elements of the inner block of the density matrix (3) through the parameters of Hamiltonian:
@@
$u_{k}=-\frac{2g_{k}}{ZR_{k}}\sinh{\frac{R_{k}}{2T}}e^{-(h_{k}+h_{2S+k})/2T}$
with $k=1,\ldots,2S$. In turn, these expressions open up an alternative opportunity to find the eigenvalues of the Gibbs density matrix using Eq. (9).
V.2 Hamiltonian and interaction constants
Expanding the Hermitian matrix (37) into generalized Gell-Mann matrices and then using the connections of these matrices with the spin ones, we obtain an expression for the Hamiltonian with $2(4S+1)$ terms of the spin matrices, their powers and products. Instead of this formal approach, one can also construct the same form of the Hamiltonian by adding linear independent spin terms with axial symmetry. In both cases, we come to the representation of the Hamiltonian (37) in the form
${\cal H}=B_{1}s_{z}+B_{2}S_{z}+J(s_{x}S_{x}+s_{y}S_{y})+J_{z}s_{z}S_{z}+D_{z}(s_{x}S_{y}-s_{y}S_{x})+K_{1}S_{z}^{2}+K_{2}s_{z}S_{z}^{2}+\ldots.$ (44)
We use a notation for the expanding coefficients which is suitable for the magnetic compounds, and will below restrict ourselves to seven parameters, namely, $B_{1}$ and $B_{2}$ are the $z$-components of an external magnetic field applied to spins 1/2 and $S$, respectively; $J$ and $J_{z}$ are the exchange Heisenberg constants; $K_{1}$ is the uniaxial one-ion anisotropy; $K_{2}$ is the uniaxial two-ion anisotropy; $D_{z}$ is the $z$-component of Dzyaloshinskii vector. It is interesting to note that the antisymmetric Dzyaloshinskii-Moriya (DM) interaction satisfies axial symmetry, while the symmetric combination $s_{x}S_{y}+s_{y}S_{x}$, which corresponds to the so-called Kaplan–Shekhtman–Entin-Wohlman–Aharony (KSEA) interaction, does not satisfy. Finally, if necessary, the Hamiltonian (44) can be extended by adding new axially invariant terms.
In Eq. (44), $s_{\alpha}=\sigma_{\alpha}/2$, where $\sigma_{\alpha}$ ($\alpha=x,y,z$) are again the Pauli matrices, and (setting $\hbar=1$)
$S_{z}={\rm diag}\,[S,S-1,\ldots,-S],$ (45)
$S_{x}=\left(\begin{array}[]{ccccccc}&b_{S}&&&&&\\ b_{S}&&b_{S-1}&&&&\\ &b_{S-1}&&b_{S-2}&&&\\ &&&&&&\\ &&&&&b_{-S+1}\\ &&&&b_{-S+1}&&\end{array}\right),$ (46)
$S_{y}=\left(\begin{array}[]{ccccccc}&-ib_{S}&&&&&\\ ib_{S}&&-ib_{S-1}&&&&\\ &ib_{S-1}&&-ib_{S-2}&&&\\ &&&&&&\\ &&&&&-ib_{-S+1}\\ &&&&ib_{-S+1}&&\end{array}\right)$ (47)
with $b_{m}=\sqrt{(S+m)(S-m+1)}/2$.
Inserting the Pauli and spin matrices (45) - (47) into Eq. (44) and then comparing the resulting open form of (44) with the matrix (37), we get expressions for the quantities $E_{0,4S+1}$, $h_{k,k+2S}$ and $g_{k}$ ($k=1,\dots,2S$) through the input parameters of the model, for each desired spin value $S$. For example, if $S=1$, then
$E_{0,5}=K_{1}+\frac{J_{z}+K_{2}}{2}\pm\Big{(}\frac{B_{1}}{2}+B_{2}\Big{)},\quad h_{1,4}=\pm\frac{B_{1}}{2},\quad h_{2,3}=-\frac{J_{z}}{2}+K_{1}\pm\Big{(}\frac{B_{1}+K_{2}}{2}-B_{2}\Big{)},\quad g_{1}=g_{2}=\frac{J+iD_{z}}{\sqrt{2}}.$ (48)
Now, one can get statistical weights (41) and entries (43), that opens up the possibility of calculating branches (24), (25), (35), (IV.2) and, finally, quantum correlations $\cal U$ and $\cal F$ through the interaction constants of the Hamiltonian (44).
VI Behavior of quantum correlations
Let us move on to consider the properties of spin-$(1/2,S)$ systems with a Hamiltonian (44). We will limit ourselves to seven specified interaction parameters because they are most important for magnetic materials.
VI.1 High-temperature wings of quantum correlations
Begin our analysis with the behavior of quantum correlations at high temperatures $T\to\infty$. Using symbolic calculations on a digital computer, we extract high-temperature asymptotics from the above general formulas. The results are as follows.
@@ VI.1 High-temperature wings of quantum correlations
${\cal F}_{0}=2\,{\cal U}_{0}=\frac{J^{2}+D_{z}^{2}}{3T^{2}}+O(1/T^{3}),\quad{\cal F}_{1}=2\,{\cal U}_{1}=\frac{3B_{1}^{2}+2K_{2}^{2}+2(J^{2}+J_{z}^{2}+D_{z}^{2}+2B_{1}K_{2})}{12T^{2}}+O(1/T^{3}).$ (49)
@@
${\cal F}_{0}=2\,{\cal U}_{0}=\frac{5(J^{2}+D_{z}^{2})}{8T^{2}}+O(1/T^{3}),\quad{\cal F}_{1}=2\,{\cal U}_{1}=\frac{16B_{1}^{2}+41K_{2}^{2}+20(J^{2}+J_{z}^{2}+D_{z}^{2}+2B_{1}K_{2})}{64T^{2}}+O(1/T^{3}).$ (50)
@@
${\cal F}_{0}=2\,{\cal U}_{0}=\frac{J^{2}+D_{z}^{2}}{T^{2}}+O(1/T^{3}),\quad{\cal F}_{1}=2\,{\cal U}_{1}=\frac{5B_{1}^{2}+34K_{2}^{2}+10(J^{2}+J_{z}^{2}+D_{z}^{2}+2B_{1}K_{2})}{20T^{2}}+O(1/T^{3}).$ (51)
@@
${\cal F}_{0}=2\,{\cal U}_{0}=\frac{35(J^{2}+D_{z}^{2})}{24T^{2}}+O(1/T^{3}),\quad{\cal F}_{1}=2\,{\cal U}_{1}=\frac{48B_{1}^{2}+707K_{2}^{2}+140(J^{2}+J_{z}^{2}+D_{z}^{2}+2B_{1}K_{2})}{192T^{2}}+O(1/T^{3}).$ (52)
These expressions show that the quantum correlations asymptotically fall to zero according to the power law $T^{-2}$. The field $B_{2}$ and strength of uniaxial single-ion anisotropy $K_{1}$ are absent here. The amplitudes of these high-temperature wings (coefficients in the $1/T^{2}$ law) demonstrate that the DM coupling $D_{z}$ and the Heisenberg interactions $J$ and $J_{z}$ enhance the quantum correlations, but the strength of these correlations decreases when the sign of the $B_{1}K_{2}$ product is negative. Note that the main high-$T$ terms ${\cal F}_{0,1}$ are twice as large as the corresponding terms ${\cal U}_{0,1}$.
VI.2 $S$ as a quantum correlation enhancer
($\rm\bf a$)Refer to caption($\rm\bf b$)
Figure 1: LQFI versus reduced temperature for the isotropic Heisenberg systems $(1/2,S)$ with $S=1$ (blue), $S=3/2$ (green), $S=2$ (orange) and $S=5/2$ (red). ($\rm\bf a$$J>0$, antiferromagnetic interactions; ($\rm\bf b$$J<0$, ferromagnetic interactions. Here, ${\cal H}=J{\vec{s}}\cdot{\vec{S}}$.
($\rm\bf a$)Refer to caption($\rm\bf b$)
Figure 2: LQFI versus reduced temperature for the isotropic Heisenberg systems $(1/2,S)$ with $S=1$ (blue), $S=3/2$ (green), $S=2$ (orange) and $S=5/2$ (red). ($\rm\bf a$$J>0$, antiferromagnetic interactions; ($\rm\bf b$$J<0$, ferromagnetic interactions. Here, ${\cal H}=\frac{J}{S}{\vec{s}}\cdot{\vec{S}}$.
Table 1: Values of ground-state quantum correlations LQU and LQFI in the spin-$(1/2,S)$ XXX dimers with both antiferromagnetic ($J>0$) and ferromagnetic ($J<0$) couplings.
$S$ $J>0$ $J<0$
1/2 1 1/3=0.333…
1 8/9=0.888… 4/9=0.444…
3/2 5/6=0.833… 1/2=0.5
2 4/5=0.8 8/15=0.533…
5/2 7/9=0.777… 5/9=0.555…
Now let us explore the behavior of quantum correlations throughout temperature region. We start with the simplest case of the completely isotropic Heisenberg XXX model setting $J_{z}=J$ and $B_{1}=B_{2}=K_{1}=K_{2}=D_{z}=0$, i.e.
${\cal H}=J{\bf s}\cdot{\bf S}.$ (53)
Calculations show that here ${\cal F}_{1}={\cal F}_{0}={\cal F}$ and ${\cal U}_{1}={\cal U}_{0}={\cal U}$. Using Eq. (35), we find LQFI for different values of spin $S$. So, when $S=1$, the quantum correlation $\cal F$ is equal to
${\cal F}^{(1)}(T)=\frac{16}{9}\frac{\sinh{(3J/4T)\tanh{(3J/4T)}}}{2+\exp{(3J/2T)}}e^{3J/4T}.$ (54)
In analogous manner, we obtain
${\cal F}^{(3/2)}(T)=\frac{5\sinh^{2}{(J/T)}}{4+4\cosh{(2J/T)}-\sinh{(2J/T)}},$ (55)
${\cal F}^{(2)}(T)=\frac{32\sinh^{2}{(5J/4T)}}{5[5+5\cosh{(5J/2T)}-\sinh{(5J/2T)}]},$ (56)
${\cal F}^{(5/2)}(T)=\frac{35}{9}\frac{\cosh{(3J/T)}-1}{6+6\cosh{(3J/T)}-\sinh{(3J/T)}}$ (57)
@@ VI.2 $S$ as a quantum correlation enhancer
Figure 1 shows the behavior of the quantum correlation $\cal F$ as a function of the reduced temperature $T/J$ (or $T/|J|$) for both antiferromagnetic and ferromagnetic systems with $S=1,\dots,5/2$. Notice, ${\cal F}$ and ${\cal U}$ coincide at $T=0$. As seen from Fig. 1, the quantum correlations have finite values at zero temperature. With increasing the temperature, the correlations first maintain a quasi-stationary value and then monotonically tend to zero.
In the case of antiferromagnetic interaction, the ground-state quantum correlations decrease with increasing spin $S$ (see Fig. 1a and Table 1). A similar picture is observed for the entanglement [53, 73]. This was interpreted as follows [53]. At $S\to\infty$, the system becomes more and more classical and therefore the quantum correlations gradually weaken and completely disappear in the limit of infinitely large spin $S$.
This interpretation is intuitively attractive, but unfortunately, it fails in the case of ferromagnets. Really, as can be seen from Fig. 1b and Table 1 (column $J<0$), quantum correlations at zero temperature surprisingly increase as spin $S$ increases. This is due to the peculiarities of the ground state.
[…]
It is interesting to draw a parallel with the behavior of quantum entanglement in the XXX model (53). Using negativity [15] as a measure of entanglement, the authors [53] found that the threshold temperature is equal to
$T_{\rm th}=\frac{2S+1}{2\ln{(2S+2)}}.$ (58)
Above this temperature, quantum entanglement vanishes. Note that quantum entanglement is present only in the case of antiferromagnetic interactions ($J>0$) and is completely absent for ferromagnets ($J<0$). From (58), a striking result follows [53]
$\lim_{S\to\infty}T_{\rm th}=\infty.$ (59)
In other words, the threshold temperature may be arbitrarily high when $S$ is sufficiently large.
However, this conclusion is controversial. Indeed, before passing to the limit $S\to\infty$, it is necessary to renormalize the spin operators: $S_{\alpha}\to S_{\alpha}/\sqrt{S(S+1)}$ ($\alpha=x,y,z$) [74, 75]. Because of this renormalization, the Hamiltonian (53) takes the form
${\cal H}=\frac{J}{\sqrt{S(S+1)}}\,{\bf s}\cdot{\bf S}\simeq\frac{J}{S}\,{\bf s}\cdot{\bf S}$ (60)
and the threshold temperature (58) should be replaced by
$T_{\rm th}=\frac{2S+1}{2S\ln{(2S+2)}}.$ (61)
From here we come to another result, namely,
$\lim_{S\to\infty}T_{\rm th}=0.$ (62)
So, increasing the spin length $S$ reduces the threshold temperature and turns it to zero in the limit of an infinitely large spin $S$.
The renormalized Hamiltonian (60) changes the behavior of quantum correlations. The new curves are shown in Fig. 2 for $S>1/2$, i.e. for relatively large spins. The main difference is occurred in the high-temperature region. Now an increase in the spin $S$ here leads to a decrease in quantum correlations. This is entirely consistent with intuitive expectations.
It is important to emphasize the following. The low-temperature limit of quantum correlations does not depend on the value of the interaction constant in the Hamiltonian (depends only on the sign of $J$). This is clearly visible from Figs. 1 and 2. Therefore, the values presented in Table 1 are universal.
VI.3 Sadden change phenomena
@@
Quantum correlations LQU and LQFI are determined by Eqs. (23) and (34), respectively. In general, these functions are piecewise-defined. The minimization conditions included in their definitions can lead to an interesting phenomenon, namely, the behavior of quantum correlations can undergo sharp changes as the parameters of the system change smoothly. In the mathematical literature, such phenomena are called catastrophes [76, 77].
@@
Consider the system, in which $B_{1}=K_{2}=D_{z}=0$. As follows from Eqs. (49) - (52), 0-wings are below 1-wings if $|J_{z}|>1$, and, conversely, above if $|J_{z}|<1$.
@@
[…]
Figure 4: Quantum correlations $\cal F$ and $\cal U$ as functions of temperature $T$ for the model with $B_{1}=0$, $B_{2}=-0.8$, $J=1$, $J_{z}=0.3$, $K_{1}=0$, $K_{2}=0$, and $D_{z}=0$ at different values of spin length $S$.
The situation begins to change dramatically at $B_{2}\neq 0$. Figure 3 shows the temperature dependencies of branches for different spin values $S$. The intersections of branches are clearly observed here. The ${\cal U}_{0}(T)$ curves intersects the ${\cal U}_{1}(T)$ ones at the points with temperatures 0.215, 0.262, 0.292 and 0.323 for $S=1$, 3/2, 2 and 5/2, respectively. These temperatures increase with increasing $S$. Similarly for the $\cal F$-branches.
On the correlation curves, the above intersection points are visible as fractures (sharp bends), see Fig. 4. They are points of non-analyticity. The functions ${\cal U}(T)$ and ${\cal F}(T)$ are continuous, but their first derivatives experience discontinuities of the first kind and therefore these functions belong to the differentiability class $C^{0}$.
@@
Figure 5: Branches of quantum correlations depending on temperature for the model with $B_{1}=-0.65$, $B_{2}=0.9$, $J=-1.3$, $J_{z}=-1$, $K_{1}=0.4$, $K_{2}=-1.02$ and $D_{z}=-0.3$ at spin lengths $S=1$ and 3/2.
As follows from Figs. 3 and 4, the above systems contain one sudden transition. It is remarkable that a series of such transitions is possible. Indeed, let $B_{1}=-0.65$, $B_{2}=0.9$, $J=-1.3$, $J_{z}=-1$, $K_{1}=0.4$, $K_{2}=-1.02$ and $D_{z}=-0.3$. As can be seen from Fig. 5, the $\cal U$-branches (blue curves) intersect twice and hence the behavior of LQU suddenly changes at two different temperatures. It is interesting to note that the second correlation, LQFI, preserves continuity and smoothness. This indicates a peculiar hybrid behavior of the non-classical correlation in the given system.
@@
Figure 6: Branches of quantum correlations depending on temperature for the model with $B_{1}=-0.65$, $B_{2}=0.9$, $J=-1.3$, $J_{z}=-1$, $K_{1}=0.4$, $K_{2}=-0.6$ and $D_{z}=-0.3$ at spin lengths $S=1$ and 3/2. Insets show the vicinities of intersection points at higher temperatures.

### v2

@@ IV System at thermal equilibrium
This matrix contains $2(4S+1)$ real free parameters (unlike density matrices, since there is no normalization condition here). Importantly, the same form is preserved for the powers of $\cal H$, their sums and hence functions (representable by Tailor series).
Expanding the Hermitian matrix (26) into generalized Gell-Mann matrices and then using the connections of these matrices with the spin ones, we obtain an expression for the Hamiltonian with $2(4S+1)$ terms of the spin matrices. Instead of this formal approach, one can also construct the same Hamiltonian by adding linear independent spin terms with axial symmetry. In both cases, we come to the representation of the Hamiltonian (26) in the Bloch-type form. In particular, the Hamiltonian of the system (1/2,$S$) with spin $S=1$ includes ten terms and can be written as
${\cal H}$ $=$ $B_{1}s_{z}+B_{2}S_{z}+J(s_{x}S_{x}+s_{y}S_{y})+J_{z}s_{z}S_{z}+KS_{z}^{2}+K_{1}(S_{x}^{2}+S_{y}^{2})+K_{2}s_{z}S_{z}^{2}+D_{z}(s_{x}S_{y}-s_{y}S_{x})$ (27)
$+$ $\Gamma[s_{x}(S_{x}S_{z}+S_{z}S_{x})+s_{y}(S_{y}S_{z}+S_{z}S_{y})]+\Lambda[s_{x}(S_{y}S_{z}+S_{z}S_{y})-s_{y}(S_{x}S_{z}+S_{z}S_{x})],$
where $s_{\mu}=\sigma_{\mu}/2$, in which $\sigma_{\mu}$ are again the Pauli matrices, and $S_{x}$, $S_{y}$ and $S_{z}$ are the components of spin $S$ (setting $\hbar=1$).
In the Hamiltonian (27), $B_{1}$ and $B_{2}$ are the $z$-components of an external magnetic field applied to spins 1/2 and 1, respectively; $J$ and $J_{z}$ are the exchange Heisenberg constants; $K$ is the single-ion uniaxial anisotropy; $K_{1}$ is the planar single-ion anisotropy; $K_{2}$ is the uniaxial two-ion anisotropy; $D_{z}$ is the $z$-component of Dzyaloshinskii vector; $\Gamma$ and $\Lambda$ are respectively symmetric and asymmetric higher-order spin coupling terms. It is interesting to note that the antisymmetric DM interaction satisfies axial symmetry, while the symmetric combination $s_{x}S_{y}+s_{y}S_{x}$, which corresponds to the so-called KSEA interaction [62], does not satisfy. The number of terms in the Hamiltonian (27) for systems with $S>1$ increases, but we will limit ourselves to the same ten terms in order to make comparative study between cases with different spins $S$.
The Gibbs density matrix is defined as
$\rho=\frac{1}{Z}\exp(-{\cal H}/T),$ (28)
where $Z={\rm Tr}\exp(-{\cal H}/T)$ is the partition function and $T$ is the temperature. Since $\cal H$ has the axial symmetrical form, this functional relationship guarantees the axial symmetry form for the matrix $\rho$.
To find Gibbs density matrix, we first rewrite the Hamiltonian (27) as an open matrix. This allows us to express all the entries of the matrix (26) through the parameters of this Hamiltonian. Next, it is clear that the energy levels of the system (26) are $E_{0}$, $E_{4S+1}$ and $2S$ additional pairs
@@
$R_{k}=\sqrt{(h_{k}-h_{2S+k})^{2}+4|g_{k}|^{2}}.$ (30)
The partition function $Z=\sum_{n}\exp(-E_{n}/T)$ is given by
@@
Finally, if the dimension of the matrix $\cal H$ is finite, $\rho$ can be found manually or obtained on a digital machine by direct symbolic (analytical) calculations. After this, the results can be extended by induction to systems with arbitrary spin $S$. This allows one to get all nonzero matrix elements of the inner block of the density matrix (2):
@@
$u_{k}=-\frac{2g_{k}}{ZR_{k}}\sinh{\frac{R_{k}}{2T}}e^{-(h_{k}+h_{2S+k})/2T}.$
So, the Gibbs density matrix has been found in terms of interaction constants of the Hamiltonian (27).
We can now move on to studying the various properties of quantum correlations.
V Behavior of quantum correlations
V.1 High-temperature region
Let us begin our analysis with the behavior of quantum correlations at high temperatures $T\to\infty$. Using formulas (13), (14), (24) and (25), we extract high-temperature asymptotics of the branches of quantum correlations for the system (27). The results are as follows.
@@ V.1 High-temperature region
${\cal F}_{0}=2\,{\cal U}_{0}=\frac{J^{2}+D_{z}^{2}+\Gamma^{2}+\Lambda^{2}}{3T^{2}}+O(1/T^{3}),\quad$ (34)
${\cal F}_{1}=2\,{\cal U}_{1}=\frac{3B_{1}^{2}+2K_{2}^{2}+2(J^{2}+J_{z}^{2}+D_{z}^{2}+\Gamma^{2}+\Lambda^{2}+2B_{1}K_{2})}{12T^{2}}+O(1/T^{3}).$ (35)
@@
${\cal F}_{0}=2\,{\cal U}_{0}=\frac{5(J^{2}+D_{z}^{2})+12(\Gamma^{2}+\Lambda^{2})}{8T^{2}}+O(1/T^{3}),\quad$ (36)
${\cal F}_{1}=2\,{\cal U}_{1}=\frac{16B_{1}^{2}+41K_{2}^{2}+20(J^{2}+J_{z}^{2}+D_{z}^{2}+2B_{1}K_{2})+48(\Gamma^{2}+\Lambda^{2})}{64T^{2}}+O(1/T^{3}).$ (37)
@@
${\cal F}_{0}=2\,{\cal U}_{0}=\frac{2(J^{2}+D_{z}^{2})+3(\Gamma^{2}+\Lambda^{2})}{2T^{2}}+O(1/T^{3}),\quad$ (38)
${\cal F}_{1}=2\,{\cal U}_{1}=\frac{5B_{1}^{2}+34K_{2}^{2}+10(J^{2}+J_{z}^{2}+D_{z}^{2}+2B_{1}K_{2})+42(\Gamma^{2}+\Lambda^{2})}{20T^{2}}+O(1/T^{3}).$ (39)
@@
${\cal F}_{0}=2\,{\cal U}_{0}=\frac{35(J^{2}+D_{z}^{2})+224(\Gamma^{2}+\Lambda^{2})}{24T^{2}}+O(1/T^{3}),\quad$ (40)
${\cal F}_{1}=2\,{\cal U}_{1}=\frac{48B_{1}^{2}+707K_{2}^{2}+140(J^{2}+J_{z}^{2}+D_{z}^{2}+2B_{1}K_{2})+896(\Gamma^{2}+\Lambda^{2})}{192T^{2}}+O(1/T^{3}).$ (41)
These expressions show that the high-temperature tails (wings) of quantum correlations monotonically tend to zero according to the power law $T^{-2}$. Thus, temperature has a destructive effect on quantum correlations in this region and completely destroys them in the limit $T\to\infty$.
Next, the $B_{2}$ field and the single-ion anisotropy strengths $K$ and $K_{1}$ are absent here; they manifest themselves in higher terms of high-temperature expansions. We also see that the main high-$T$ terms ${\cal F}_{0,1}$ are exactly twice as large as the corresponding terms of ${\cal U}_{0,1}$.
Equations (34) - (41) make it easy to obtain valuable information, namely 0- or/and 1-branches determine the behavior of quantum correlations ${\cal U}=\min{\{{\cal U}_{0},{\cal U}_{1}\}}$ and ${\cal F}=\min{\{{\cal F}_{0},{\cal F}_{1}\}}$ in the high-temperature region.
All these features will be important in the following subsections when analyzing the behavior of quantum correlations over the entire temperature region.
V.2 Increasing quantum correlations with temperature
The parameter space of the studied model (27) is $R^{10}$. The space is huge and there is no possibility to present all the obtained graphic material. Therefore, we will limit ourselves to only some discovered features of quantum correlations.
($\rm\bf a$)Refer to caption($\rm\bf b$)Refer to caption($\rm\bf c$)
Figure 1: Quantum correlations $\cal F$ and $\cal U$ as functions of temperature $T$ for the qubit-qutrit model with $B_{1}=B_{2}=K=K_{1}=K_{2}=D_{z}=0$, $J=1$, $J_{z}=-1.7$, $\Lambda=1$, and $\Gamma=1$ (a), 0.38 (b) and 0.3 (c).
@@ V.2 Increasing quantum correlations with temperature
The behavior at high temperatures has been established above. More interesting features and peculiarities of quantum correlations are observed in lower temperature regions.
Let us consider the XXZ model with only $\Gamma$- and $\Lambda$-interactions; $J=1$, $J_{z}=-1.7$, $\Lambda=1$, and $\Gamma$ is a variable parameter. Look at Fig. 1. From Fig. 1a it is clearly seen that at $\Gamma=1$ both correlations LQU and LQFI monotonically tend to zero with increasing temperature. Thus, as the system heats up, quantum correlations gradually weaken. This is in line with our intuitive expectations.
As the interaction $\Gamma$ decreases, monotonicity breaks down, as can be seen from Fig. 1b, where $\Gamma=0.38$. The curves are deformed, minima appear near $T\approx 0.076$ and maxima at $T\approx 0.29$ for $\cal U$ and at $T\approx 0.47$ for $\cal F$. It is important to emphasize that both quantum correlations increase with an increase in temperature $T$ in the local interval between these minima and maxima. And this happens despite the fact that, generally speaking, temperature causes degradation of the coherence and quantum nature of physical systems.
Moreover, temperature can generate nonclassical correlations near the ground state and then increase them. This is clearly shown by the curves in Fig. 1c, which corresponds to $\Gamma=0.3$. In the temperature range from zero to $T\approx 0.32$, LQU increases from zero to about 0.31, and LQFI increases from zero to about 0.34 when the temperature varies from zero to $T\approx 0.48$.
Figure 3: Quantum correlations $\cal F$ and $\cal U$ as functions of temperature $T$ for the model with $B_{1}=0$, $B_{2}=-0.8$, $J=1$, $J_{z}=0.3$ and $K=K_{1}=K_{2}=D_{z}=\Gamma=\Lambda=0$ at different values of spin length $S$.
Thus, temperature destroys the quantum correlations of a physical system at high temperatures. On the other hand, heating the system can lead to a local increase in its quantumness in the ground state and above it.
Finally, we note that the behavior of LQU and LOFI qualitatively reproduces each other in all cases presented in Fig. 1.
V.3 Sadden change phenomena
@@
Quantum correlations LQU and LQFI are determined by Eqs. (12) and (23), respectively. In general, these functions are piecewise-defined. The minimization conditions included in their definitions can lead to an interesting phenomenon, namely, the behavior of quantum correlations can undergo sharp changes as the parameters of the system change smoothly. In the mathematical literature, such phenomena are called catastrophes [63].
@@
Consider the system, in which $B_{1}=K=K_{1}=K_{2}=D_{z}=\Gamma=\Lambda=0$, i.e. there is XXZ dimer in an external magnetic field $B_{2}$. As follows from Eqs. (34) - (41), 0-branches are below 1-branches if $|J_{z}|>1$, and, conversely, above if $|J_{z}|<1$.
@@
The situation begins to change dramatically at $B_{2}\neq 0$. Figure 2 shows the temperature dependencies of branches for different spin values $S$. The intersections of branches are clearly observed here. The ${\cal U}_{0}(T)$ curves intersect the ${\cal U}_{1}(T)$ ones at the points with temperatures 0.215, 0.262, 0.292 and 0.323 for $S=1$, 3/2, 2 and 5/2, respectively. These temperatures increase with increasing $S$. Similarly, they are true for the $\cal F$-branches except that the intersects occur at different temperatures.
On the correlation curves, the above intersection points are visible as fractures (sharp bends), see Fig. 3. They are non-analytical points. The functions ${\cal U}(T)$ and ${\cal F}(T)$ are continuous, but their first derivatives experience discontinuities of the first kind and therefore these functions belong to the differentiability class $C^{0}$.
@@
Figure 4: Branches of quantum correlations depending on temperature for the model with $B_{1}=-0.65$, $B_{2}=0.9$, $J=-1.3$, $J_{z}=-1$, $K=0.4$, $K_{2}=-1.02$, $D_{z}=-0.3$ and $K_{1}=\Gamma=\Lambda=0$ at spin lengths $S=1$ and 3/2.
@@
Figure 5: Branches of quantum correlations depending on temperature for the model with $B_{1}=-0.65$, $B_{2}=0.9$, $J=-1.3$, $J_{z}=-1$, $K=0.4$, $K_{2}=-0.6$, $D_{z}=-0.3$ and $K_{1}=\Gamma=\Lambda=0$ at spin lengths $S=1$ and 3/2. Insets show the vicinities of intersection points at higher temperatures.
As follows from Figs. 2 and 3, the above systems contain one sudden transition. It is remarkable that a series of such transitions is possible. Indeed, let $B_{1}=-0.65$, $B_{2}=0.9$, $J=-1.3$, $J_{z}=-1$, $K=0.4$, $K_{2}=-1.02$, $D_{z}=-0.3$ and $K_{1}=\Gamma=\Lambda=0$. As can be seen from Fig. 4, the $\cal U$-branches (blue curves) intersect twice and hence the behavior of LQU suddenly changes at two different temperatures. It is interesting to note that the second correlation, LQFI, preserves continuity and smoothness. This indicates a peculiar hybrid behavior of the non-classical correlation in the given system.